# On the Natural Robustness of Vision-Language Models Against Visual Perception Attacks in Autonomous Driving

### v2

@@ Abstract
Autonomous vehicles rely on deep neural networks for critical tasks such as traffic sign recognition, automated lane centering, and vehicle detection. However, these models are vulnerable to attacks that can cause misclassifications and compromise safety. Defense mechanisms, including adversarial training, often degrade benign accuracy and fail to generalize against unseen attacks. In this work, we introduce Vehicle Vision Language Models (V2LMs), fine-tuned Vision-Language Models (VLMs) specialized for AV perception. Our findings demonstrate that V2LMs inherently exhibit superior robustness against unseen attacks without requiring adversarial training, maintaining significantly higher accuracy than conventional DNNs under adversarial conditions. We evaluate two deployment strategies: Solo Mode, where individual V2LMs handle specific perception tasks, and Tandem Mode, where a single unified V2LM is fine-tuned for multiple tasks simultaneously. Experimental results reveal that DNNs suffer performance drops of 33%–46% under attacks, whereas V2LMs maintain adversarial accuracy with reductions of less than 8% on average. The Tandem Mode further offers a memory-efficient alternative while achieving comparable robustness to the Solo Mode. Also, we explore the integration of V2LMs as parallel components to AV perception to enhance resilience against adversarial threats. Our results suggest that V2LMs offer a promising path toward more secure and resilient AV perception systems. [^1] The implementation code of this work is available at https://github.com/pedram-mohajer/V2LM
1 Introduction
Autonomous vehicles (AVs) are built on automated driving systems (ADS), consisting of modular subsystems including perception, planning, and control. The perception module—significantly enhanced by advancements in deep neural network (DNN) models for tasks such as image classification and object detection [62, 40]—reads and processes data from sensors such as cameras and LiDAR [41, 44]; this information is then sent to the planning module, which makes decisions on navigation and maneuvers, and subsequently to the control module, executing these decisions to control the vehicle’s movement [52].
While these DNN models effectively recognize, monitor, and predict the movements of nearby objects [55, 70], they are also vulnerable to physical attacks that manipulate real-world environments to deceive perception systems [72, 24, 38, 39]. Such physical attacks pose a significant threat to AV perception by causing misclassification and unsafe driving behaviors, as demonstrated by methods such as DRP-Attack [51], ControlLoc [35], and SlowTrack [34].
@@ 1 Introduction
1(a) shows that adversarial training (AT) reduces YOLOv5-cls’s accuracy on benign inputs by over 20%, while NVILA-8B and Qwen-VL-7B remain stable, with changes under 3%—suggesting that, unlike conventional models, VLMs are less sensitive to the trade-offs typically introduced by AT [46]. 1(b) shows that both VLMs also achieve high accuracy on adversarial inputs, with NVILA-8B staying near 98% and Qwen-VL-7B near 85%. In contrast, YOLOv5-cls starts at 55.8% under attack and gains only 13.2% with AT [61].
@@
To address these limitations and achieve both robust and generalizable models for AV perception tasks, this work proposes a novel finding: fine-tuned VLMs inherently exhibit robustness against unseen adversarial attacks on AV perception systems even without adversarial training, significantly outperforming DNNs. As shown in 1(b), VLMs such as Qwen-VL [4] and NVILA [32] demonstrate strong resilience, achieving 82.93% and 98.04% accuracy respectively against adversarial examples without adversarial training. In contrast, the YOLOv5-cls [26, 47] model suffers a drastic drop from 94.88% accuracy on benign inputs to just 55.82% on adversarial ones. Based on this observation, we introduce fine-tuned VLMs for the first time as a robust AV perception module, naming them Vehicle Vision Language Models (V2LMs) and comprehensively evaluating them in the AV context.
@@
This paper evaluates six VLMs, namely LLaVA-7B, LLaVA-13B-LoRA [30], MoE-LLaVA [28], MobileVLM [15], Qwen-VL-7B, and NVILA-8B for their potential as auxiliary in-vehicle components to enhance perception performance in the presence of attacks targeting perception algorithms. The first four models are based on LLaMA [59], while the last two—Qwen-VL-7B and NVILA-8B—are based on Qwen [3], selected for their state-of-the-art performance in multimodal tasks and their efficiency in offline operation—an essential feature for our application. In addition to these VLMs, the study also evaluates task-specific DNN models, including YOLOv5-cls [26] for Traffic Sign Recognition (TSR), CLRerNet [21] for Automated Lane Centering (ALC), and YOLOv5-dt [47] for Vehicle Detection (VD), as strong baselines under both benign and adversarial conditions.
Before ATAfter AT000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111AccuracyNVILAQwen-VLYOLOv5-cls
(a) Benign
Before ATAfter AT000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Accuracy
(b) Adversarial
Figure 1: Impact of adversarial training (AT) on NVILA, Qwen-VL, and YOLOv5-cls. (a) Benign inputs: AT reduces clean accuracy for YOLOv5-cls while VLMs remain high. (b) Adversarial inputs: VLMs (NVILA/Qwen-VL) exhibit strong robustness even before adversarial training; YOLOv5-cls still lags.
The study begins by evaluating the VLMs’ zero-shot performance on key AV tasks, including (i) TSR, (ii) ALC, and (iii) VD as part of Object Detection (OD). Subsequently, they are fine-tuned specifically for these tasks to better align them with AV applications, with their performance reassessed to determine improvements (RQ1). We refer to each fine-tuned model as a Vehicle Vision-Language Model (V2LM), highlighting its specialization for AV tasks. Additionally, the task-specific DNN-based models are fine-tuned on the same training dataset as the VLMs, ensuring a fair comparison of their capabilities under identical conditions.
@@
Then, the effectiveness and resilience of both DNN models and V2LMs are tested using unseen adversarial examples (AEs) against three distinct attacks targeting AV perception algorithms (RQ2): (1) Robust and Accurate UV-map-based Camouflage attack (RAUCA) to deceive VD algorithms [73], (2) a physical-world adversarial attack known as the Dirty Road Patch (DRP-Attack) to compromise DNN-based automated lane centering (ALC) models [51], and (3) shadows cast on traffic signs to attack TSR algorithms [72].
@@
The study then compares two distinct designs for utilizing V2LMs and evaluates their performance on the three aforementioned AV tasks and AEs (RQ3). The first design, termed Solo Mode, involves separate V2LMs, each fine-tuned individually for one of the AV tasks. The second design, named Tandem Mode, uses a single V2LM fine-tuned simultaneously for all three AV tasks, aiming to improve robustness against various attack types. By comparing these two designs, we aim to determine whether a unified model can achieve robustness across multiple tasks and adversarial conditions as effectively as task-specific models.
Finally, we discuss the feasibility of integrating a V2LM as a parallel component within AV systems to bolster the perception module against attacks. Given the critical need for low latency to ensure real-time decision-making in autonomous driving, V2LM’s latency is evaluated relative to the perception system and human reaction times. This assessment aims to determine whether V2LM can operate within strict real-time constraints without adversely impacting the vehicle’s responsiveness.
This paper makes the following contributions:
@@
  • This work proposes a novel finding that fine-tuned Vision Language Models (VLMs) inherently exhibit superior robustness against unseen adversarial attacks compared to task-specific DNNs. We highlight a critical limitation of defense methods, such as adversarial training: while adversarial training aims to improve robustness, it significantly degrades the benign accuracy of DNNs while providing only limited improvements against adversarial examples. In contrast, VLM-based models achieve strong adversarial robustness while maintaining high benign accuracy.
@@
  • We introduce Vehicle Vision Language Models (V2LMs)—fine-tuned VLMs specifically for AV perception tasks: TSR, ALC, and VD. We further propose two deployment strategies: Solo Mode, where separate VLM are fine-tuned individually for each task, and Tandem Mode, employing a single unified V2LM across multiple perception tasks in same time.
@@
  • We conduct comprehensive experiments to evaluate the robustness of DNN models and V2LMs under adversarial conditions. DNN models experience performance drops of 33%–46% under attacks, whereas V2LMs achieve significantly smaller reductions (less than 8% on average), consistently maintaining high adversarial accuracy without relying on additional defense mechanisms.
Table 1: Zero-shot performance of VLMs on AV tasks (Solo Design).
Task Model Accuracy F1-Score Precision Recall
TSR LLaVA-7B 8.23% 9.19% 2.38% 2.19%
LLaVA-13B-LoRA 13.51% 15.03% 4.93% 6.04%
MoE-LLaVA 1.19% 1.63% 2.06% 0.48%
MobileVLM 1.27% 0.95% 1.15% 0.81%
Qwen-7B 7.01% 1.73% 4.52% 6.08%
NVILA-8B 31.24% 29.80% 29.24% 30.31%
ALC LLaVA-7B 37.16% 36.56% 42.71% 29.33%
LLaVA-13B-LoRA 38.20% 38.12% 45.86% 25.64%
MoE-LLaVA 35.41% 28.84% 29.68% 21.39%
MobileVLM 20.71% 24.60% 28.38% 21.71%
Qwen-7B 26.96% 11.67% 7.45% 26.97%
NVILA-8B 50.91% 51.18% 51.44% 50.91%
VD LLaVA-7B 77.91% 64.01% 88.09% 51.61%
LLaVA-13B-LoRA 91.26% 84.46% 89.82% 80.23%
MoE-LLaVA 75.61% 65.16% 87.95% 50.46%
MobileVLM 60.64% 67.83% 71.43% 64.57%
Qwen-7B 75.44% 82.57% 94.38% 75.44%
NVILA-8B 92.06% 92.01% 91.82% 92.06%
2 Related Work
Large Language Models in Autonomous Driving.
@@ 2 Related Work
Recent work has demonstrated the potential of LLMs in the context of AVs, particularly enhancing perception, control, and motion planning tasks. Regarding perception systems, LLMs utilize external APIs to access real-time information sources, including traffic reports, and weather updates, which significantly enrich the vehicle’s ability to gain a comprehensive understanding of its environment [16]. Aldeen et al. [2] investigate the application of Large Multimodal Models (LMMs) for enhancing the cybersecurity of AVs. Concerning control, LLMs enable the adjustment of control settings according to driver preferences, thereby personalizing the driving experience [53].
Additionally, LLMs enhance transparency by providing detailed explanations of each step in the control process. These capabilities extend to navigation improvements; e.g., LLMs effectively process real-time traffic data to identify congested routes and suggest alternative paths, significantly optimizing navigation for both efficiency and safety [56]. For motion planning tasks, LLMs leverage their advanced natural language understanding and reasoning capabilities [37]. This allows for enhanced user-centric communication, enabling passengers to articulate their intentions and preferences in everyday language. Furthermore, LLMs analyze textual data from sources such as maps, traffic reports, and real-time updates to make high-level decisions that optimize route planning [42].
Adversarial Threats and Defenses.
Adversarial Examples (AEs) were first defined by Szegedy et al. [58] as subtly modified inputs designed to fool DNNs. These minor, often imperceptible alterations can drastically alter a DNN’s predictions [54, 9]. Several studies have demonstrated the vulnerability of AV perception systems to AEs. Kong et al. [27] introduced PhysGAN, a GAN-based framework that generates AEs resilient in the physical world, capable of misleading ADS throughout an entire trajectory. Similarly, Eykholt et al. [18] showed that even placing a sticker on a stop sign could deceive the AV’s TSR.
To counter these vulnerabilities, various defense mechanisms have been proposed. Adversarial training, proposed by Madry et al. [36], involves training models on both clean and AEs to enhance robustness. Defensive distillation, introduced by Papernot et al. [43], aims to reduce the model’s sensitivity to perturbations by smoothing decision boundaries. Robust optimization techniques, discussed by Wong and Kolter [63], improve the model’s resistance to adversarial attacks through advanced optimization methods. Also, input preprocessing techniques, such as those outlined by Guo et al. [20], involve denoising or transforming inputs to mitigate the effects of adversarial perturbations.
3 V2LM as a Defense Mechanism
3.1 Overview
@@
The perception system in AVs is responsible for interpreting the environment through sensor data, such as images, to enable safe and efficient operation. This system performs essential tasks including TSR, ALC, and and VD which is part of object detection [68]. TSR enables the vehicle to follow road rules by recognizing traffic signs [60], ALC keeps the vehicle centered within its lane by identifying road markings [17], and VD detects and classifies other vehicles [8, 11]. The perception system (PS) can be represented as a function, shown in Equation 1, which takes an input image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$ and outputs the results of TSR, ALC, and VD. Specifically, TSR outputs the detected traffic sign’s class and bounding box if a sign is present; ALC provides a classification for the appropriate steering command; and VD returns the bounding box and class of any detected car.
@@
where $f$ denotes the specific target module in the AV’s perception system, which can be the TSR, ALC, or VD module. This perturbation $\delta$, constrained by $\epsilon$, represents a small, controlled modification designed to be imperceptible, thereby ensuring that $\mathcal{I}_{\text{adv}}$ visually resembles $\mathcal{I}$. Despite this, the adversarial perturbation may lead to misclassification of traffic signs, incorrect lane-following commands, or faulty detection of objects.
3.2 Vision Language Model (VLM)
VLMs, by integrating visual and textual data [69], have the potential to enhance perception tasks by enabling a more comprehensive understanding of the environment. A pre-trained VLM takes an image and a text prompt as inputs to generate a relevant response:
[…]
@@ 3.2 Vision Language Model (VLM)
3.3 Vehicle Vision Language Model (V2LM)
Fine-tuning VLMs on AV-specific tasks is essential for optimizing their performance and enabling adaptation to domain-specific challenges such as variations in lighting, road conditions, and traffic scenarios; the resulting fine-tuned VLM is referred to as V2LM. For LLaVA, this fine-tuning process involves adjustments to key components. The vision encoder, implemented as CLIP ViT-L/14 [45], extracts visual features from input images. The language model, based on Vicuna (a fine-tuned variant of LLaMA) [14], interprets textual prompts. A cross-modal attention module integrates the modalities, and all components are jointly optimized via visual instruction tuning [31] to align with autonomous vehicle (AV) tasks.
@@ 3.3 Vehicle Vision Language Model (V2LM)
LoRA (Low-Rank Adaptation) [22] offers an efficient alternative by focusing on specific parameters within the model. Only low-rank matrices are introduced in certain layers — primarily within the self-attention and feedforward blocks — allowing the majority of the pre-trained parameters to remain frozen. This technique reduces the memory and computational load while achieving task-specific adjustments by updating only the new, smaller matrices. By leveraging fine-tuning techniques for AV tasks and the efficiency gains of LoRA, exploring V2LMs as a solution to AV perception challenges has the potential to address critical limitations of existing methods, enhancing robustness against adversarial attacks while avoiding the performance degradation common in traditional defenses.
V2LM ${}^{TSR}_{\text{Solo}}$2LM ${}^{ALC}_{\text{Solo}}$2LM ${}^{VD}_{\text{Solo}}$romptTSRDataTSROutputTSRPromptALCDataALCOutputALCPromptVDDataVDOutputVD
@@
Figure 2: Comparison of Solo and Tandem Modes
Solo vs. Tandem Design Comparison.
Deploying V2LMs to enhance robustness against AEs in AVs raises important considerations for their implementation. Given the diverse range of tasks that AVs must perform, it is crucial to determine the best strategy for utilizing them. One approach involves using a separate V2LM for each specific AV task to improve robustness within each module, ensuring specialized and precise detection capabilities. This design, referred to as Solo Mode, is illustrated in 2(a), where individual V2LMs are dedicated to tasks such as TSR, ALC, and VD. The formal representation of this design is:
[…]
Alternatively, a single V2LM can handle all AV tasks in a tandem approach, providing a unified method for improving robustness across multiple modules, as depicted in 2(b). In this design, multiple image-query pairs—each corresponding to a different task—are combined into a single input, where images are concatenated using a separator, and queries are merged in the same order. The model processes this structured batch simultaneously within one forward pass, extracting task-specific outputs for each image-query pair independently while treating the collection as a cohesive input during execution:
$\{\textit{Output}_{\textit{TSR}},\textit{Output}_{\textit{ALC}},\textit{Output}_{\textit{VD}}\}={}$ (5)
[…]
@@
V2LM-Augmented Perception Systems.
Integrating V2LMs with AV perception systems offers the potential to strengthen resilience against AEs in real-time. Currently, AV perception systems face significant limitations in mitigating attacks, which can lead to dangerous misinterpretations of sensor data. As Cao et al. [10] demonstrated, despite efforts to enhance the security of AV perception, significant vulnerabilities persist and traditional detection mechanisms often fail to mitigate these threats, leading to potentially dangerous consequences for AV decision-making and safety.
By integrating V2LM in the end-to-end AV stack, AVs can benefit from its ability to process data, thereby working in parallel with perception tasks to enhance robustness and support accurate decision-making. The outputs of the perception system and V2LM can then be used by the downstream planning and control modules to act accordingly in the presence of AEs. This context is depicted in Figure 3.
It is crucial to evaluate the latency of V2LM integration to ensure it can enhance robustness against AEs in real-time without affecting the efficiency of the perception system. AV vision algorithms can process images at a rate of 2-4 frames per second (fps), which corresponds to approximately 250-500 ms per image [6]. As a result, V2LM should process inputs within the required real-time constraints to support robust perception.
4 Experimental Evaluation
4.1 Experimental Setup
To evaluate the efficacy of V2LMs, the focus was on three critical tasks: Traffic Sign Recognition (TSR), Automated Lane Centering (ALC) and Vehicle Detection (VD). For the former, the German Traffic Sign Recognition Benchmark (GTSRB) [57] was utilized, which includes images captured under various conditions such as different lighting and distances. Each picture from this dataset, which has 42 classes, was sent with the prompt ”Identify this traffic sign.” to the respective V2LM. For the ALC task, a dataset was generated using the CARLA simulator [12], which includes 6,000 training and 2,000 testing images. These images, taken from the driver’s perspective under various weather conditions and times of day, are classified into three categories indicating the next move: Straight, Left, and Right. The prompt used for this task was, ”As a car driver, at which direction should you turn the steering wheel?”

### v3

# Toward Inherently Robust VLMs Against Visual Perception Attacks

@@ Abstract
Autonomous vehicles rely on deep neural networks (DNNs) for traffic sign recognition, lane centering, and vehicle detection, yet these models are vulnerable to attacks that induce misclassification and threaten safety. Existing defenses (e.g., adversarial training) often fail to generalize and degrade clean accuracy. We introduce Vehicle Vision–Language Models (V2LMs), fine-tuned Vision Language Models (VLMs) specialized for AV perception, and show that they are inherently more robust to unseen attacks without adversarial training, maintaining substantially higher adversarial accuracy than conventional DNNs. We study two deployments: Solo (task-specific V2LMs) and Tandem (a single V2LM for all three tasks). Under attacks, DNNs drop 33%–74%, whereas V2LMs decline by under 8% on average. Tandem achieves comparable robustness to Solo while being more memory-efficient. We also explore integrating V2LMs in parallel with existing perception stacks to enhance resilience. Our results suggest V2LMs are a promising path toward secure, robust AV perception. Code and data are available at https://github.com/pedram-mohajer/V2LM.
I Introduction
Autonomous vehicles (AVs) rely on automated driving systems (ADS) consisting of perception, planning, and control. The perception module uses deep neural network (DNN) models for image classification and object detection, processes camera and LiDAR data [47], and passes the resulting scene understanding to the planning module for navigation decisions, which are then executed by the control module [38]. Prior point-cloud analysis methods include plane-kernel convolutions [31] and non-parametric designs using Gaussian positional encoding [salarpour2024pointgn]. Although these DNN-based perception models recognize, monitor, and predict nearby objects [54], they are vulnerable to physical attacks that manipulate real-world environments to deceive perception [20], causing misclassification and unsafe driving behaviors, as shown by the DRP-Attack [37] on lane detection. Similar reliability concerns also arise across other applied ML pipelines beyond autonomous driving, including medical imaging and analytics and scientific simulation workflows [41, 36, 1].
@@ I Introduction
Before ATAfter AT00.20.20.40.40.60.60.80.811AccuracyNVILAQwen-VLYOLOv5-clsViT-small
(a) Accuracy on adversarial
Before ATAfter AT00.20.20.40.40.60.60.80.811Accuracy
(b) Accuracy on benign
Figure 1: Impact of adversarial training (AT) on NVILA, Qwen-VL, YOLOv5-cls, and ViT-small. (a) Adversarial inputs. (b) Benign inputs.
@@
To address these limitations and achieve robust and generalizable models for AV perception tasks, this work proposes a novel finding: VLMs inherently exhibit robustness against unseen adversarial attacks on AV perception systems even without adversarial training, significantly outperforming DNNs. VLMs such as Qwen-VL [4] and NVILA [27] demonstrate strong resilience against adversarial inputs without adversarial training. Based on this observation, we are the first to systematically fine-tune VLMs for AV perception tasks; fine-tuning trains these VLMs on TSR, ALC, and VD, enabling a fair comparison with task-specific DNNs, and we name them Vehicle Vision Language Models (V2LMs).
Thus, we benchmark both conventional DNNs and V2LMs before and after adversarial training (AT) on clean and adversarial inputs under unseen adversarial attacks (“unseen” means these attacks are never used during training or fine-tuning of the DNNs or VLMs and appear only at test time). As shown in Figure 1, this comparison highlights two findings. (i) VLMs exhibit robustness: before AT, NVILA-8B remains near 98% and Qwen-VL about 83% on adversarial inputs, while both stay above 95% on benign inputs; after AT, these numbers change by no more than $\pm$3 percentage points (pp), underscoring their insensitivity to the usual AT trade-offs [33]. (ii) Conventional DNNs collapse under attack, and AT provides only modest relief—and at a cost: YOLOv5-cls [22] plunges from roughly 95% to 56% (-39 pp), and ViT-small [14] drops from 96% to 23% (-73 pp); although AT increases adversarial accuracy to 69% for YOLOv5-cls and 60% for ViT-small, their benign accuracies simultaneously sink to 74% and 68%, reflecting the clean-data penalty noted in previous work [46].
TABLE I: Classification accuracy (%) of YOLOv5-cls and ViT-small on GTSRB without defenses (row Base Model) and accuracy change $\Delta$ ($p$) when standard defenses are applied to benign and Shadow images. Green ($\uparrow$) and red ($\downarrow$) entries denote improvement or degradation vs. the no-defense Base Model.
YOLOv5-cls ViT-small
Method benign shadow benign shadow
Base Model 94.88 55.82 96.20 23.12
Label Smoothing (LS) $\uparrow 0.87$ $\downarrow 3.71$ $\downarrow 2.47$ $\downarrow 11.71$
Dropout (DO) $\downarrow 3.17$ $\downarrow 8.26$ $\downarrow 1.85$ $\downarrow 10.06$
JPEG Compression (JPG) $\downarrow 2.81$ $\downarrow 11.79$ $\downarrow 1.74$ $\downarrow 8.03$
Bit-Depth Reduction (BD) $\downarrow 1.75$ $\downarrow 0.95$ $\downarrow 0.53$ $\uparrow 2.27$
Random Resized Cropping (RRP) $\downarrow 22.00$ $\downarrow 25.42$ $\downarrow 4.22$ $\downarrow 14.53$
Histogram Equalization (HEQ) $\uparrow 1.59$ $\downarrow 0.82$ $\downarrow 1.03$ $\uparrow 1.12$
@@
We evaluate six VLMs (LLaVA-7B, MobileVLM [12], LLaVA-13B-LoRA [25], MoE-LLaVA [23], Qwen-VL-7B [4], and NVILA-8B [27]) as auxiliary perception modules for traffic-sign recognition (TSR), automated lane centering (ALC), and vehicle detection (VD) tasks. Each model is tested zero-shot and then re-evaluated after task-specific fine-tuning (RQ1). The first four models are LLaMA-based, the last two Qwen-based, chosen for their performance and offline suitability. Task-specific DNN baselines such as YOLOv5-cls and ViT-small for TSR, CLRerNet for ALC, and YOLOv5-dt for VD are fine-tuned on the same data to ensure a fair comparison with the VLMs under both benign and adversarial conditions.
@@
Then, the effectiveness and resilience of both DNN models and V2LMs are tested using unseen (not used during model training or fine-tuning) adversarial examples (AEs) against three distinct attacks targeting AV perception algorithms (RQ2): (1) Robust and Accurate UV-map-based Camouflage attack (RAUCA) to deceive VD algorithms [57], (2) a physical-world adversarial attack known as the Dirty Road Patch (DRP-Attack) to compromise DNN-based automated lane centering (ALC) models, and (3) shadows cast on traffic signs to attack TSR algorithms [56]; these are strong physical attacks evaluated end-to-end and shown to cause safety-relevant failures.
@@
The study then compares two distinct designs for utilizing V2LMs and evaluates their performance on the three aforementioned AV tasks and AEs (RQ3). The first design, termed Solo Mode, involves separate V2LMs, each fine-tuned individually for one of the AV tasks. The second design, named Tandem Mode, uses a single V2LM fine-tuned simultaneously for all three AV tasks, aiming to assess whether it can match the robustness of Solo Mode across tasks under both benign and adversarial conditions. This paper makes the following contributions:
@@
  • This work presents a novel finding: fine-tuned VLMs inherently exhibit superior robustness against unseen adversarial attacks compared to task-specific DNNs. We highlight a critical limitation of adversarial training: it significantly degrades benign accuracy while providing only limited improvements against adversarial examples. In contrast, VLMs achieve strong adversarial robustness while maintaining high benign accuracy.
@@
  • We introduce Vehicle Vision Language Models (V2LMs) which are fine-tuned VLMs for AV perception tasks: TSR, ALC, and VD. We propose two deployment strategies: Solo Mode (separate VLM fine-tuned for each task) and Tandem Mode (a single unified V2LM across multiple perception tasks).
@@
  • We conduct comprehensive experiments to evaluate the robustness of DNN models and V2LMs under adversarial conditions. DNN models experience performance drops of 33%–74% under attacks, whereas V2LMs show reductions of less than 8% on average, maintaining high adversarial accuracy without additional defense mechanisms.
II Related Work
@@
Large Language Models in Autonomous Driving. Recent works have demonstrated the potential of LLMs in the context of AVs, particularly enhancing perception, control, and motion planning tasks. Regarding perception systems, LLMs utilize external APIs to access real-time information sources, including traffic reports, and weather updates, which significantly enrich the vehicle’s ability to gain a comprehensive understanding of its environment [13]. Aldeen et al. [2] investigate the application of Large Multimodal Models (LMMs) for enhancing the cybersecurity of AVs. Concerning control, LLMs enable the adjustment of control settings according to driver preferences, thereby personalizing the driving experience [39].
Adversarial Threats and Defenses. Adversarial Examples (AEs) were first defined by Szegedy et al. [44] as subtly modified inputs designed to fool DNNs. These minor, often imperceptible alterations can drastically alter a DNN’s predictions [40, 7]. To counter these vulnerabilities, various defense mechanisms have been proposed. We re-implemented these methods and evaluated them against the shadow attack [56]; as summarized in Table I, they are largely ineffective and often reduce benign accuracy. Label Smoothing (LS) [43] softens targets and can underfit fine-grained TSR cues. Dropout (DO) [42] removes capacity needed for small text and borders. JPEG Compression (JPG) [17] suppresses high-frequency edges and characters. Bit-Depth Reduction (BD) [50] distorts class-specific hues via quantization. Random Resized Cropping (RRP) [49] can crop out discriminative regions and destabilize scale. Histogram Equalization (HEQ) [58] over-equalizes, amplifying halos and shifting the distribution.
III V2LM as a Defense Mechanism
III-A Autonomous Vehicle Perception
@@
The perception system in AVs is responsible for interpreting the environment through sensor data, such as images, to enable safe and efficient operation. This system performs essential tasks including TSR, ALC, and and VD which is part of object detection. TSR enables the vehicle to follow road rules by recognizing traffic signs [45], ALC keeps the vehicle centered within its lane by identifying road markings [15], and VD detects and classifies other vehicles [9]. The perception system (PS) can be represented as a function, shown in Equation 1, which takes an input image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$ and outputs the results of TSR, ALC, and VD. Specifically, TSR outputs the detected traffic sign’s class and bounding box if a sign is present; ALC provides a classification for the appropriate steering command; and VD returns the bounding box and class of any detected car.
Refer to captionRefer to caption
(a) Shadow Attack
(b) DRP-Attack
(c) RAUCA
Figure 2: Examples of adversarial attacks targeting AV perception.
@@
where $f$ denotes the specific target module in the AV’s perception system, which can be the TSR, ALC, or VD module. This perturbation $\delta$, constrained by $\epsilon$, represents a small, controlled modification designed to be imperceptible, thereby ensuring that $\mathcal{I}_{\text{adv}}$ visually resembles $\mathcal{I}$.
III-B Vision Language Model (VLM)
Definition. VLMs, by integrating visual and textual data [53], have the potential to enhance perception tasks by enabling a more comprehensive understanding of the environment. A pre-trained VLM takes an image and a text prompt as inputs to generate a relevant response:
$\textit{VLM}_{\textit{pre}}(\mathcal{I},\textit{Prompt})\rightarrow\textit{Generated-Response}$ (3)
@@ III-B Vision Language Model (VLM)
Fine-Tuning. Fine-tuning VLMs on AV-specific tasks is essential for optimizing their performance and enabling adaptation to domain-specific challenges such as variations in lighting, road conditions, and traffic scenarios; the resulting fine-tuned VLM is referred to as V2LM. For the LLaVA family of models, this fine-tuning process involves adjustments to key components. The vision encoder, implemented as CLIP ViT-L/14 [32], extracts visual features from input images. The language model, based on Vicuna (a fine-tuned variant of LLaMA) [11], interprets textual prompts. A cross-modal attention module integrates the modalities, and all components are jointly optimized via visual instruction tuning [26] to align with autonomous vehicle (AV) tasks.
@@
V2LM Perception System Iroadview+Outputall
LoRA (Low-Rank Adaptation) [18] offers an efficient alternative by focusing on specific parameters within the model. Only low-rank matrices are introduced in certain layers, primarily within the self-attention and feedforward blocks, allowing the majority of the pre-trained parameters to remain frozen. This technique reduces the memory and computational load while achieving task-specific adjustments by updating only the new, smaller matrices. By leveraging fine-tuning techniques for AV tasks and the efficiency gains of LoRA, exploring V2LMs as a solution to AV perception challenges has the potential to address critical limitations of existing methods, enhancing robustness against adversarial attacks while avoiding the performance degradation common in traditional defenses.
V2LM $^{\text{TSR}}_{\text{Solo}}$V2LM $^{\text{ALC}}_{\text{Solo}}$V2LM $^{\text{VD}}_{\text{Solo}}$PromptTSRDataTSROutputTSRPromptALCDataALCOutputALCPromptVDDataVDOutputVD
@@
Figure 4: Comparison of Solo and Tandem Modes.
Solo vs. Tandem Design Comparison. Deploying V2LMs to enhance robustness against AEs in AVs raises important considerations for their implementation. Given the diverse range of tasks that AVs must perform, it is crucial to determine the best strategy for utilizing them. One approach involves using a separate V2LM for each specific AV task to improve robustness within each module, ensuring specialized and precise detection capabilities. This design, referred to as Solo Mode, is illustrated in 4a, where individual V2LMs are dedicated to tasks such as TSR, ALC, and VD. The formal representation of this design is:
$\textit{Output}_{i}={\textit{V\textsuperscript{2}LM}}^{i}_{\textit{solo}}(\textit{Data}_{i},\textit{Prompt}_{i})$ (4)
where $i\in\{\textit{TSR},\textit{ALC},\textit{VD}\}$. Here, $\textit{Data}_{i}$ denotes the dataset specific to each perception task $T_{i}$, and $\textit{Prompt}_{i}$ is the prompt customized to fine-tune ${\textit{V\textsuperscript{2}LM}}^{i}_{\textit{solo}}$ for that particular task, ensuring task-specific optimization.
Alternatively, a single V2LM can handle all AV tasks using a tandem approach, providing a unified method to improve robustness across multiple modules, as depicted in 4b. In this design, multiple image query pairs, each corresponding to a different task, are combined into a single input: the images are concatenated using a separator and the queries are merged in the same order. The model processes this structured batch simultaneously within one forward pass, extracting task-specific outputs for each image-query pair independently while treating the collection as a cohesive input during execution:
$\{\textit{Output}_{\textit{TSR}},\textit{Output}_{\textit{ALC}},\textit{Output}_{\textit{VD}}\}={\textit{V\textsuperscript{2}LM}}_{\textit{tandem}}(\textit{Data}_{\textit{all}},\textit{Prompt}_{\textit{all}})$
@@
TABLE II: Range of zero-shot performance (%) across these six VLMs for each AV perception task. Results reflect the lowest and highest values among all models before fine-tuning.
Task Accuracy (%) F1-Score (%) Precision (%) Recall (%)
TSR 1.19–31.24 0.95–29.80 1.15–29.24 0.48–30.31
ALC 20.71–50.91 24.60–51.18 28.38–51.44 21.39–50.91
VD 60.64–92.06 64.01–92.01 71.43–94.38 50.46–92.06
Integrating V2LMs with AV perception systems offers the potential to strengthen resilience against AEs in real-time. Currently, AV perception systems face significant limitations in mitigating attacks, which can lead to dangerous misinterpretations of sensor data. As Cao et al. [8] demonstrated, despite efforts to enhance the security of AV perception, significant vulnerabilities persist and traditional detection mechanisms often fail to mitigate these threats, leading to potentially dangerous consequences for AV decision-making and safety. By integrating V2LM in the end-to-end AV stack, AVs can benefit from its ability to process data, thereby working in parallel with perception tasks to enhance robustness and support accurate decision-making. The outputs of the perception system and V2LM can then be used by the downstream planning and control modules to act accordingly in the presence of AEs. This context is depicted in Figure 3.
It is crucial to evaluate the latency of V2LM integration to ensure it can enhance robustness against AEs in real-time without affecting the efficiency of the perception system. AV vision algorithms can process images at a rate of 2-4 frames per second (fps), which corresponds to approximately 250-500 ms per image [6].
IV Experimental Evaluation
IV-A Experimental Setup
To evaluate the efficacy of V2LMs, the focus was on three critical tasks: Traffic Sign Recognition (TSR), Automated Lane Centering (ALC) and Vehicle Detection (VD). For the former, the German Traffic Sign Recognition Benchmark (GTSRB) was utilized, which includes images captured under various conditions such as different lighting and distances. Each picture from this dataset, which has 42 classes, was sent with the prompt ”Identify this traffic sign.” to the respective V2LM. For the ALC task, a dataset was generated using the CARLA simulator, which includes 6,000 training and 2,000 testing images. These images, taken from the driver’s perspective under various weather conditions and times of day, are classified into three categories indicating the next move: Straight, Left, and Right. The prompt used for this task was, ”As a car driver, at which direction should you turn the steering wheel?”